Solve by elimination
-4x-13= 5y
7x+5y= -19

Found 2 solutions by Cromlix, FrankM:
Answer by Cromlix(4381)

Hi there,
-4x-13= 5y
7x+5y= -19
Rearrange
-4x - 5y = 13......(1)
7x + 5y = -19......(2)
Add (1) + (2)
3x = -6
x = -2
Substitute x = -2 into (2)
7x + 5y = -19
7(-2) + 5y = -19
-14 + 5y = -19
5y = -19 + 14
5y = -5
y = -1
{-2,-1}
Hope this helps :-)

Answer by FrankM(1040)

-4x-13= 5y
7x+5y= -19
rearrange eq1
-4x-5y=13
7x+5y= -19
3x=-6
x=-2
-4(-2)-5y=13
8-5y=13
-5y=5
y=-1
(-2,-1)